# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v2

@@ A.5 Photons $(\gamma)$
Photons are massless and therefore ultra-relativistic. Unlike non-relativistic particles, one must account for the direction $\cos\theta=\boldsymbol{p}\cdot\boldsymbol{k}/\lvert\boldsymbol{p}\rvert\lvert\boldsymbol{k}\rvert$ of their momenta $\boldsymbol{p}$ relative to the Fourier wavenumber $\boldsymbol{k}$. This results in a theoretically infinite hierarchy of equations for Legendre multipoles $l$, which in practice must be truncated at some maximum multipole $l_{\text{max}}$:
@@
The equations for $F_{l}^{\prime}$ and $G_{l}^{\prime}$ apply for $2\leq l<l_{\text{max}}$. There are no tight-coupling, radiation-streaming or ultra-relativistic fluid approximations. Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e. $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e. $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=-\frac{8}{15}\frac{k}{\kappa^{\prime}}F_{1}$, $G_{0}=\frac{5}{16}F_{2}$, $G_{1}=-\frac{1}{16}\frac{k}{\kappa^{\prime}}F_{2}$, $G_{2}=\frac{1}{16}F_{2}$, and $F_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}F_{l-1}$ and $G_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}G_{l-1}$ for $3\leq l\leq l_{\text{max}}$. The species is parametrized by its temperature $T_{0}$ today, which in turn sets the density parameters $\Omega_{0}=\frac{\pi^{2}}{15}\frac{(k_{B}T_{0})^{4}}{(\hbar c)^{3}}\frac{8\pi G}{3H_{0}^{2}}$ and $\rho_{0}=\frac{8\pi}{3}\Omega_{0}$ today.
@@ A.6 Massless neutrinos $(\nu)$
Massless neutrinos behave similarly to photons, but decouple from interactions with other species in the very early universe. One must only account for this interaction in initial conditions, while their evolution equations are a simpler case of the photons’:
@@
The equations for $F_{l}^{\prime}$ apply for $2\leq l<l_{\text{max}}$. There is no ultra-relativistic fluid approximation. Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e. $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e. $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=\frac{2}{15}(k\tau)^{2}\Psi$ and $F_{l}=\frac{l}{2l+1}k\tau F_{l-1}$. The species is parametrized by the effective number $N_{\text{eff}}$, the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ and temperature $T_{0}$ today. If photons are present, they default to $T_{\nu 0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ and $\Omega_{\nu 0}=N_{\text{eff}}\frac{7}{8}\big(\frac{4}{11}\big)^{4/3}\Omega_{\gamma 0}$.
@@ A.7 Massive neutrinos $(h)$
Massive neutrinos are the most complicated species in the $\mathrm{\Lambda CDM}$ model (alongside baryon recombination). In essence, the species we have looked at so far have Boltzmann equations where the momenta of their distribution function can be integrated out analytically in non-relativistic and ultra-relativistic limits. This means that their stress-energy components are linked by trivial equations of state and sound speeds, for example, and their effect can be parametrized by a simple density parameter $\Omega_{0}$.
On the other hand, massive neutrinos have intermediate masses that fall between the non-relativistic and ultra-relativistic limits. Integrals over their distribution function must be computed numerically. This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that minimizes the number of sampled points. Fortunately, the momentum integrals have a structure that can be exploited: they are all in the form weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma and Bertschinger (1995) for more details). One can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations). In other words, the integral operator $\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)$ is effectively replaced by the discrete summation operator $\sum_{i}\!W_{i}$ for some weights $W_{i}$.
On top of this, perturbations are also expanded in Legendre multipoles $l$ up to a cutoff $l_{\text{max}}$:
@@
Initial conditions are $\psi_{i,0}=-\frac{1}{4}(-2\Psi)\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$, $\psi_{i,1}=-\frac{1}{3}\frac{E_{i}}{x_{i}}\frac{1}{2}k\tau\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$, $\psi_{i,2}=-\frac{1}{2}\frac{1}{15}(k\tau)^{2}\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$ and $\psi_{i,l}=0$. This integrates to adiabatic $\delta/(1+w)$, $\theta$ and $\sigma$ similarly to massless neutrinos. Free parameters are the temperature today $T_{0}$, mass $m$ of a single neutrino and degeneracy factor $N=\sum_{i=1}^{N}m_{i}/m$ for describing multiple neutrinos with equal mass. The degeneracy factor defaults to $N=3$, and the temperature to $T_{h0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ if photons are present, as for massless neutrinos.
Here the equation for $\psi_{i,l}^{\prime}$ applies for $2\leq l\leq l_{\text{max}}$, and expressions $g_{i}=g(x_{i})$ indexed by $i$ are evaluated with the momentum quadrature point $x=x_{i}$. The reduction to dimensionless momenta $x=pc/k_{B}T$ (the argument of $\exp$ in $f$) is deliberate because it makes numerics more well-defined and the quadrature scheme independent of $m$ and all other cosmological parameters.
@@
This substitution achieves two things: the scaling $x/L$ brings the dominant integral contributions well within $x/L\ll 1$ if $L$ is chosen to be a characteristic decay ?length? of the distribution function, and the rational part $1/(1+x/L)$ maps the infinite domain $x\in(0,\infty)$ to the finite domain $u\in(0,1)$, which can be integrated numerically. The substituted integrand is then passed to an adaptive algorithm in QuadGK.jl [^18] https://github.com/JuliaMath/QuadGK.jl that computes quadrature points $u_{i}$ and weights $W_{i}$ by performing weighted integrals against several test functions $g(x)$. Finally, the corresponding momenta $x_{i}=x(u_{i})$ are returned along with the weights $W_{i}$, from which one can approximate the integral $I[g(x)]\approx\sum_{i}W_{i}g(x_{i})$ against any $g(x)$.
SymBoltz tests this numerical quadrature scheme against the analytical result $I[x^{n-2}]=\int_{0}^{\infty}\mathrm{d}x\,x^{n}/(e^{x}+1)=(1-2^{-n})\zeta(n+1)\Gamma(n+1)$ for $2\leq n\leq 8$. We assume this to be a reasonable test for the integrals encountered in the equations above. Agreement is excellent with $L=100$, which yields relative errors below $10^{-6+n-N}$ for all $2\leq n\leq 8$ and $1\leq N\leq 5$. SymBoltz defaults to $N=4$ momenta, for which this relative error is less than $10^{-6}$ for $n\leq 4$, for example. It also agrees well with CLASS using default settings.
Note that this momentum quadrature strategy is generic with respect to the distribution function $f(x)$ and substitution $u(x)$, so it can easily be modified for other particle species whose distribution function cannot be integrated out.
CAMB (Lewis, 2025) and CLASS (Lesgourgues and Tram, 2011) apply similar weighted quadrature strategies. They also get away with only a handful of sampled momenta, but the precise details of the computation differ slightly. For reference, here are points and weights computed by SymBoltz for $1\leq N\leq 8$ momenta:
@@ A.8 Cosmological constant $(\Lambda)$
It is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today. This is set to $\Omega_{\Lambda 0}=1-\sum_{s\neq\Lambda}\Omega_{s0}$ if all species $s$ have a $\Omega_{s0}$ parameter and General Relativity is the theory of gravity. This constraint comes from the 1st Friedmann equation today.
@@ A.9 Primordial power spectrum ($I$)
SymBoltz computes the inflationary primordial power spectrum parametrized by the amplitude $A_{s}$ and tilt $n_{s}$:
@@ A.10 Matter power spectrum
SymBoltz computes the matter power spectrum for some desired set of species $s$, which are presumably matter-like at late times (e.g. $s\in\{c,b,h\}$):
@@
Here $\Delta$ is the total gauge-independent overdensity with total $\delta$ and $\theta$ computed by summing the components of the energy-momentum tensor that are additive.
@@ A.11 CMB power spectrum and line-of-sight integration
SymBoltz finds photon temperature and polarization multipoles today for any $l$ by computing the line-of-sight integrals
@@
As first suggested by Seljak and Zaldarriaga (1996), this approach enables cheap computation for any $l$ after integrating the perturbation ODEs with only a few $l\leq l_{\text{max}}$. This drastically speeds up the computation over including all $l$ in an enormous set of coupled perturbation ODEs. SymBoltz performs the integrals with the trapezoid method using the substitution $u(\tau)=\tanh(\tau)$, which adds more points in the early universe when sampled uniformly, using $768$ points by default. Here $j_{l}$ are the spherical Bessel functions of the first kind. SymBoltz is not yet generalized to non-flat geometries, where they are replaced by hyperspherical functions. The cross-correlated angular spectrum between $\mathrm{A},\mathrm{B}\in\{\mathrm{T},\mathrm{E}\}$ is then computed from
@@ A.12 Adaptive source function refinement
SymBoltz can optionally refine source functions $S(\tau,k)$ in $k$, such as when computing the matter or CMB power spectra.
First, the source function $S_{i}(\tau)=S(\tau,k_{i})$ is found on an initial grid of wavenumbers $k_{i}$ by integrating their perturbations. The method then iterates over each interval $(k_{i},k_{i+1})$, integrates the perturbations for the center $k_{i+1/2}=(k_{i}+k_{i+1})/2$ and finds $S_{i+1/2}(\tau)=S(\tau,k_{i+1/2})$. The idea is then to compare this to a linearly interpolated source $S_{\text{int}}(\tau)=(S_{i}(\tau)+S_{i+1}(\tau))/2$ with some error measure err, for example from $\texttt{err}^{2}=\int\mathrm{d}\tau\big(S_{\text{int}}(\tau)-S_{i+1/2}(\tau)\big)^{2}$. If the interpolation is poor and err is greater than some absolute and relative tolerances, the left and right half-intervals $(k_{i},k_{i+1/2})$ and $(k_{i+1/2},k_{i+1})$ are recursively refined in the same way.
This gives $S(\tau,k)$ by integrating as few kk–modes as possible. Instead of manually creating a non-uniform $k$-grid with many hyperparameters, the process is controlled by only one tolerance. It parallelizes over each interval and can also be done in $\log k$.
@@ Appendix B Precision parameters for CLASS
When comparing results to CLASS in section 3, we use CLASS version 3.3.4 with the following precision parameters: {codebox}
tight_coupling_approximation = 5 # compromise_CLASS; cannot disable TCAtight_coupling_trigger_tau_c_over_tau_h = 0.01 # turn off TCA earliertight_coupling_trigger_tau_c_over_tau_k = 0.001 # turn off TCA earlierradiation_streaming_approximation = 3 # turn off RSAncdm_fluid_approximation = 3 # turn off NCDMFAur_fluid_approximation = 3 # turn off UFAevolver = 1 # for implicit ndf15; 0 for explicit rktol_perturbations_integration = 1.0e-6 # less noise; varied in work-precision diagramtol_ncdm_newtonian = 1.0e-5 # default sampling of massive neutrino momentabackground_Nloga = 6000 # for stable derivativesWe run CLASS with output = mPk to compute only $P(k)$, and output = tCl, pCl to compute only $C_{l}^{\text{TT}}$, $C_{l}^{\text{EE}}$ and $C_{l}^{\text{TE}}$ with the code’s most specialized and optimized paths. CLASS and SymBoltz are configured to use the same Newtonian gauge, $l_{\text{max}}$ for all species, RECFAST recombination model, $\tanh$-like reionization parametrization, $w_{0}w_{a}$ dark energy equations and cosmological parameter values. The settings above disable as many approximations as possible and reduce the impact of the tight-coupling approximation, which cannot be disabled. We had to decrease the parameter background_Nloga from the default value 40000 to make the derivatives in fig. 4 stable. This parameter controls the number of points used for splining background functions in the perturbations. Its default value was changed from 3000 to 40000 in 2023, but we suspect that the increased density in points makes the splines oscillate from numerical noise. Likewise, the tolerance for the perturbations ODEs is set one notch smaller than the default $10^{-5}$ to make the output stable. We use CLASS’ default sampling of massive neutrino momenta, as discussed in section 4.
When running CLASS with all approximations on in tables 1 and 7, the first block of parameters above are not used. In the work-precision comparison in fig. 7, we also vary tol_perturbations_integration and evolver as described in the figure.

### v3

@@ A.5 Photons $(\gamma)$
Photons are massless and therefore ultrarelativistic. Unlike nonrelativistic particles, we must account for the direction $\cos\theta=\boldsymbol{p}\cdot\boldsymbol{k}\,/\,\lvert\boldsymbol{p}\rvert\cdot\lvert\boldsymbol{k}\rvert$ of their momenta $\boldsymbol{p}$ relative to the Fourier wavenumber $\boldsymbol{k}$. This results in a theoretically infinite hierarchy of equations for Legendre multipoles $l$, which in practice must be truncated at some maximum multipole $l_{\text{max}}$:
@@
The equations for $F_{l}^{\prime}$ and $G_{l}^{\prime}$ hold for $2\leq l<l_{\text{max}}$. There are no tight-coupling, radiation-streaming or ultrarelativistic fluid approximations. Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e., $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e., $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=-\frac{8}{15}\frac{k}{\kappa^{\prime}}F_{1}$, $G_{0}=\frac{5}{16}F_{2}$, $G_{1}=-\frac{1}{16}\frac{k}{\kappa^{\prime}}F_{2}$, $G_{2}=\frac{1}{16}F_{2}$, $F_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}F_{l-1}$, and $G_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}G_{l-1}$ for $3\leq l\leq l_{\text{max}}$. The species is parameterized by its temperature $T_{0}$ today, which in turn sets the density parameters $\Omega_{0}=\frac{\pi^{2}}{15}\frac{(k_{B}T_{0})^{4}}{(\hbar c)^{3}}\frac{8\pi G}{3H_{0}^{2}}$ and $\rho_{0}=\frac{8\pi}{3}\Omega_{0}$ today.
@@ A.6 Massless neutrinos $(\nu)$
Massless neutrinos behave similarly to photons, but decouple from interactions with other species in the very early universe. We must only account for this interaction in initial conditions, while their evolution equations are a simpler case of the photon equations:
@@
The equations for $F_{l}^{\prime}$ hold for $2\leq l<l_{\text{max}}$. There is no ultrarelativistic fluid approximation. Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e., $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e., $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=\frac{2}{15}(k\tau)^{2}\Psi$, and $F_{l}=\frac{l}{2l+1}k\tau F_{l-1}$. The species is parameterized by the effective number $N_{\text{eff}}$, the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today, and temperature $T_{0}$ today. If photons are present, they default to $T_{\nu 0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ and $\Omega_{\nu 0}=N_{\text{eff}}\frac{7}{8}\big(\frac{4}{11}\big)^{4/3}\Omega_{\gamma 0}$.
@@ A.7 Massive neutrinos $(h)$
Massive neutrinos are the most complicated species in the $\mathrm{\Lambda CDM}$ model (alongside baryon recombination). In essence, the species we have looked at so far have Boltzmann equations where the momenta of their distribution function can be integrated out analytically in nonrelativistic or ultrarelativistic limits. This means that their stress-energy components are linked by trivial equations of state and sound speeds, for example, and their effect can be parameterized by a simple density parameter $\Omega_{0}$.
On the other hand, massive neutrinos have intermediate masses that fall between the nonrelativistic and ultrarelativistic limits. Integrals over their distribution function must be computed numerically. This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that samples as few momenta as possible. Fortunately, the momentum integrals have a structure that can be exploited: they are all in the weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma and Bertschinger 1995 for more details). We can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations). In other words, the integral operator $\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)$ is effectively replaced by the discrete summation operator $\sum_{i}\!W_{i}$ for some weights $W_{i}$. On top of this, perturbations are also expanded in Legendre multipoles $l$ up to a cutoff $l_{\text{max}}$:
@@
Initial conditions are $\psi_{i,0}=-\frac{1}{4}(-2\Psi)\big(\frac{\mathrm{d}\ln f}{\mathrm{d}\ln x}\big)_{i}$, $\psi_{i,1}=-\frac{1}{3}\frac{E_{i}}{x_{i}}\frac{1}{2}k\tau\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d}\ln x}\big)_{i}$, $\psi_{i,2}=-\frac{1}{2}\frac{1}{15}(k\tau)^{2}\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d}\ln x}\big)_{i}$, and $\psi_{i,l}=0$. When integrated over momenta, they are equivalent to adiabatic $\delta/(1+w)$, $\theta$, and $\sigma$, similarly to massless neutrinos. Free parameters are the temperature $T_{0}$ today, single neutrino mass $m$, and degeneracy factor $N=\sum_{i=1}^{N}m_{i}/m$ for describing multiple neutrinos with the same mass. They default to $N=3$, and $T_{h0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ if photons are present, as for massless neutrinos.
Here, the equations for $\psi_{i,l}^{\prime}$ hold for $2\leq l\leq l_{\text{max}}$, and the $i$-indexed expressions $g_{i}=g(x_{i})$ are evaluated with the momentum quadrature point $x=x_{i}$. The reduction to dimensionless momenta $x=pc/k_{B}T$ (the argument of $\exp$ in $f$) is deliberate because it makes numerics more well-defined and the quadrature scheme independent of $m$ and all other cosmological parameters.
@@
This substitution achieves two things: the scaling $x/L$ brings the dominant integral contributions well within $x/L\ll 1$ if $L$ is chosen to be a characteristic decay ?length? of the distribution function, and the rational part $1/(1+x/L)$ maps the infinite domain $x\in(0,\infty)$ to the finite domain $u\in(0,1)$, which can be integrated numerically. The substituted integrand is then passed to an adaptive algorithm in QuadGK.jl [^23] https://github.com/JuliaMath/QuadGK.jl that computes quadrature points $u_{i}$ and weights $W_{i}$ by performing weighted integrals against several test functions $g(x)$. Finally, the corresponding momenta $x_{i}=x(u_{i})$ are returned along with the weights $W_{i}$, from which we can approximate the integral $I[g(x)]\approx\sum_{i}W_{i}g(x_{i})$ against any $g(x)$.
We test this numerical quadrature scheme against the analytical result $I[x^{n-2}]=\int_{0}^{\infty}\mathrm{d}x\,x^{n}/(e^{x}+1)=(1-2^{-n})\zeta(n+1)\Gamma(n+1)$ for $2\leq n\leq 8$. We assume this is a reasonable test for the integrals encountered in the equations above. The agreement is excellent with $L=100$, which yields relative errors below $10^{-6+n-N}$ for all $2\leq n\leq 8$ and $1\leq N\leq 5$. SymBoltz defaults to $N=4$ momenta, for which this relative error is less than $10^{-6}$ for $n\leq 4$, for example. It also agrees well with CLASS using default settings.
We note that this momentum quadrature strategy is generic with respect to the distribution function $f(x)$ and substitution $u(x)$, so it can easily be modified for other particle species whose distribution function cannot be integrated out.
CAMB [^24] https://cosmologist.info/notes/CAMB.pdf and CLASS (Lesgourgues and Tram 2011) apply similar weighted quadrature strategies. They also get away with only a handful of sampled momenta, but the precise details of the computation differ slightly. For reference, here are points and weights computed by SymBoltz for $1\leq N\leq 8$ momenta:
@@ A.8 Cosmological constant $(\Lambda)$
It is parameterized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today. If all species $s$ have a $\Omega_{s0}$ parameter and general relativity is the theory of gravity, it is set to $\Omega_{\Lambda 0}=1-\sum_{s\neq\Lambda}\Omega_{s0}$. This constraint comes from the first Friedmann equation today.
@@ A.9 Primordial power spectrum ($I$)
SymBoltz computes the inflationary primordial power spectrum parameterized by the amplitude $A_{s}$ and tilt $n_{s}$:
@@ A.10 Matter power spectrum
SymBoltz computes the matter power spectrum for some desired set of species $s$, which are presumably matter-like at late times (e.g., $s\in\{c,b,h\}$):
@@
Here, $\Delta$ is the total gauge-independent overdensity with total $\delta$ and $\theta$ computed by summing the components of the energy-momentum tensor that are additive.
@@ A.11 CMB power spectrum and line-of-sight integration
SymBoltz finds photon temperature and polarization multipoles today for any $l$ by computing line-of-sight integrals:
@@
As first suggested by Seljak and Zaldarriaga (1996), this approach enables cheap computation for any $l$ after integrating the perturbation ODEs with only a few $l\leq l_{\text{max}}$. This drastically speeds up the computation over including all $l$ in an enormous set of coupled perturbation ODEs. SymBoltz performs the integrals with the trapezoid method using around $1250$ uniformly sampled $\tau$, by default. Here, $j_{l}$ are the spherical Bessel functions of the first kind. SymBoltz is not yet generalized to curved spacetimes, where they are replaced by hyperspherical functions. The cross-correlated angular spectrum between $\mathrm{A},\mathrm{B}\in\{\mathrm{T},\mathrm{E}\}$ is then computed from
@@ A.12 Adaptive source function refinement
SymBoltz can optionally refine source functions $S(\tau,k)$ in $k$, such as when computing the matter or CMB power spectra. First, the source function $S_{i}(\tau)=S(\tau,k_{i})$ is found on an initial grid of wavenumbers $k_{i}$ by integrating their perturbations. The method then iterates over each interval $(k_{i},k_{i+1})$, integrates the perturbations for the center $k_{i+1/2}=(k_{i}+k_{i+1})/2$, and finds $S_{i+1/2}(\tau)=S(\tau,k_{i+1/2})$. The idea is then to compare this to a linearly interpolated source $S_{\text{int}}(\tau)=(S_{i}(\tau)+S_{i+1}(\tau))/2$ with some error measure err, for example from the $L_{2}$-like $\texttt{err}^{2}=\int\mathrm{d}\tau\big(S_{\text{int}}(\tau)-S_{i+1/2}(\tau)\big)^{2}$. If the interpolation is poor and err is greater than some absolute and relative tolerances, the left and right half-intervals $(k_{i},k_{i+1/2})$ and $(k_{i+1/2},k_{i+1})$ are recursively refined in the same way.
This gives $S(\tau,k)$ while integrating as few kk–modes as possible. Instead of manually creating a nonuniform $k$-grid with many hyperparameters, the process is controlled by only one tolerance. It parallelizes over each interval and can also be done in $\log k$.
@@ Appendix B Precision parameters for CLASS
When comparing results to CLASS in Sect. 3, we use CLASS version 3.3.4 with the following precision parameters: {codebox}
tight_coupling_approximation = 5 # compromise_CLASS; cannot disable TCAtight_coupling_trigger_tau_c_over_tau_h = 0.01 # turn off TCA earliertight_coupling_trigger_tau_c_over_tau_k = 0.001 # turn off TCA earlierradiation_streaming_approximation = 3 # turn off RSAncdm_fluid_approximation = 3 # turn off NCDMFAur_fluid_approximation = 3 # turn off UFAevolver = 1 # for implicit ndf15; 0 for explicit rktol_perturbations_integration = 1.0e-6 # less noise; varied in work-precision diagramtol_ncdm_newtonian = 1.0e-5 # default sampling of massive neutrino momentabackground_Nloga = 6000 # for stable derivativesWe run CLASS with output = mPk to compute only $P(k)$, and output = tCl, pCl to compute only $C_{l}^{\text{TT}}$, $C_{l}^{\text{EE}}$ and $C_{l}^{\text{TE}}$ with the code’s most specialized and optimized paths. CLASS and SymBoltz are configured to use the same Newtonian gauge, $l_{\text{max}}$ for all species, RECFAST recombination model, $\tanh$-like reionization parameterization, $w_{0}w_{a}$ dark energy equations, and cosmological parameter values. The settings above disable as many approximations as possible and reduce the impact of the tight-coupling approximation, which cannot be disabled. We had to decrease the parameter background_Nloga from the default value 40000 to make the derivatives in Fig. 4 stable. This parameter controls the number of points used for splining background functions in the perturbations. Its default value was changed from 3000 to 40000 in 2023, but we suspect that the increased density in points makes the splines oscillate from numerical noise. Likewise, the tolerance for the perturbations ODEs is set one notch smaller than the default $10^{-5}$ to make the output stable. We use CLASS’ default sampling of massive neutrino momenta, as discussed in Sect. 4.
When running CLASS with all approximations on in Table 1 and Fig. 7, the first block of parameters above are not used. In the work-precision comparison in Fig. 7, we also vary tol_perturbations_integration and evolver as described in the figure.